A car is traveling north along a straight road with a constant speed of 11.2 m/s. In order to pass through an intersection before the traffic light turns red, the driver accelerates forward with a constant magnitude of 2.25 m/s2 for 3.70 s. A)What is v0? B) What is a? ​C)What is t? D) What is the velocity of the car at the end of the acceleration? E)What is the position of the car at the end of the acceleration? F) How far did the car travel during the period of acceleration?
Answer #1

vo =11.2m/s, a=2.25m/s^2, t=3.7s

A) initial velocity vo =11.2m/s,

B) acceleration a =2.25m/s^2,

C) time t=3.7s

D) v =vo+a*t =11.2+2.25*3.7 =19.525m/s

E) assuming start of acceleration at origin,

s=u*t+1/2*a*t^2 =11.2*3.7+0.5*2.25*3.7^2 =(56.84 j^)m i.e along north direction

F) S=56.84m